OpenStudy (anonymous):

I need help proving this by induction. Thanks Prove that n 2 > n+ 1 for all integer n ≥ 2.

ganeshie8 (ganeshie8):

\[\large \begin{align} n \ge 2 \implies n^2 &\ge 2n \\~\\&= n+n \\~\\& \gt n+1\end{align}\]

ganeshie8 (ganeshie8):

Oh you want to prove it by induction is it

ganeshie8 (ganeshie8):

induction should be rpetty straightforward, where are you stuck at ?